# ACTIVE: Continuous Similarity Search for Vessel Trajectories

### v2

@@ Abstract.
Publicly available vessel trajectory data is emitted continuously from the global AIS system. Continuous trajectory similarity search on this data has applications in, e.g., maritime navigation and safety. Existing proposals typically assume an offline setting and focus on finding similarities between complete trajectories. Such proposals are less effective when applied to online scenarios, where similarity comparisons must be performed continuously as new trajectory data arrives and trajectories evolve. We therefore propose a real-time continuous trajectory similarity search method for vessels (ACTIVE). We introduce a novel similarity measure, object-trajectory real-time distance, that emphasizes the anticipated future movement trends of vessels, enabling more predictive and forward-looking comparisons. Next, we propose a segment-based vessel trajectory index structure that organizes historical trajectories into smaller and manageable segments, facilitating accelerated similarity computations. Leveraging this index, we propose an efficient continuous similar trajectory search (CSTS) algorithm together with a variety of search space pruning strategies that reduce unnecessary computations during the continuous similarity search, thereby further improving efficiency. Extensive experiments on two large real-world AIS datasets offer evidence that ACTIVE is capable of outperforming state-of-the-art methods considerably. ACTIVE significantly reduces index construction costs and index size while achieving a 70% reduction in terms of query time and a 60% increase in terms of hit rate.
Tiantian Liu and Hengyu Liu have equal contributions.
Tianyi Li is the corresponding author.
1. Introduction
@@
Automatic Identification System (AIS) data, which provides real-time information on vessel positions, speeds, and directions, offers a rich resource for analyzing vessel movements. With the growing availability of AIS data, understanding vessel trajectories has become essential for navigation safety (Li et al., 2024b) and environmental monitoring (Jin et al., 2021). Trajectory similarity search (Hu et al., 2023) plays a key role in applications such as collision avoidance (Li et al., 2022), route optimization (Xu et al., 2024), and movement pattern analysis (Hu et al., 2024).
@@
Given the dynamic nature of maritime environments, real-time continuous trajectory similarity search is particularly valuable. It enables timely and proactive decision-making by identifying similar movement patterns as new AIS data arrives, facilitating responsive navigation and operational safety. Unlike traditional methods, which rely on retrospective analysis of complete trajectories, real-time similarity search operates incrementally, processing incoming data without requiring an entire trajectory. Moreover, real-time search considers the future movement trends and the destination of a moving object. For instance, as illustrated in Figure 1, for a moving vessel $o$, given its current trajectory $T_{o}$ and target destination $p_{o}^{d}$, we continuously search for similar historical trajectories, accounting for both past movements and the vessel’s future trend. Despite significant progress in real-time trajectory similarity search, existing studies (Ding et al., 2008; Xie et al., 2017; Wang et al., 2018; Li et al., 2020; He et al., 2022; Luo and Zeng, 2023; Bakalov et al., 2005b, a; Ta et al., 2017; Shang et al., 2017) face limitations, leading to the following challenges.
Figure 1. Example of trajectories at timestamp $t$
@@
Challenge 1: How to effectively measure trajectory similarity. Existing methods, such as Dynamic Time Warping (Yi et al., 1998; Ying et al., 2016), Longest Common Subsequence (Vlachos et al., 2002), Fréchet distance (Alt and Godau, 1995; De Berg et al., 2013; Agarwal et al., 2014), Hausdorff distance (Hangouet, 1995; Bai et al., 2011), and Edit distance-based measures (Chen and Ng, 2004; Chen et al., 2005; Ranu et al., 2015), are commonly used to identify similar trajectories by analyzing historical movement data. However, they rely heavily on static or full-path comparisons, making them of low utility for evolving trajectories that need to be analyzed incrementally as new data arrives. Next, they focus exclusively on historical movements, overlooking the importance of an object’s near-future movement and target destination. This limits their effectiveness in real-time applications, such as collision detection, traffic management, and route prediction, where anticipating future movements is crucial. For example, in Figure 1, the most similar trajectory for $T_{o}$ is $T_{3}$ using Fréchet or Hausdorff distance. However, when considering the destination of the object, the most similar trajectory should be $T_{1}$.
Challenge 2: How to improve efficiency and scalability in continuous similarity trajectory search. Existing similar trajectory search methods (Ding et al., 2008; Xie et al., 2017; Wang et al., 2018; Li et al., 2020; He et al., 2022; Luo and Zeng, 2023) are usually designed for searches using traditional similarity metrics intended for historical data, making them less effective for real-time applications. Moreover, they are designed for offline, batch processing on static datasets, requiring full re-computations when new data arrives. This is unsuitable in a setting with continuous updates to AIS data.
@@
Achieving efficient, real-time continuous similarity searches for vessel trajectories involves addressing several interconnected challenges. First, vessel trajectories are complex, often involving irregular and dynamic movements in open water. In contrast, on road networks, structured paths and intersections simplify trajectory analysis. Second, pruning irrelevant or low-similarity candidates with high accuracy is particularly complex in a dynamic environment where trajectory data changes continuously. Effective pruning methods must strike a balance between reducing candidate sets and maintaining the accuracy of similarity computations. Third, scaling a system and retaining fast response times when handling continuous queries over extended periods adds another layer of difficulty. The system must be capable of managing large-scale, real-time data streams while maintaining high accuracy and adapting to varying query loads. Together, these challenges call for new innovations to achieve a scalable and accurate trajectory search system.
@@
Contributions. In response to these challenges, we present a novel framework for real-time continuous trajectory similarity search for vessels. To address Challenge 1, we propose a novel similarity measure called object-trajectory real-time distance, which prioritizes the future trends of vessel movements and their target destinations while accounting also for past trajectory parts and the significance of individual trajectory points. This measure enables more accurate predictions of future vessel positions and behaviors, making it better suited for real-time decision-making in maritime navigation and traffic management.
@@
To address Challenge 2, we propose a novel index structure called the segment-based vessel trajectory index, which organizes historical vessel trajectories by partitioning them into smaller, manageable segments using a Minimum Bounding Rectangle (MBR)-optimization method. This partitioning accelerates similarity computations by enabling a focus on the most relevant trajectory candidates, avoiding the need for exhaustive analyses of entire trajectories. Moreover, we design an efficient continuous similar trajectory search algorithm with four speed-up strategies that exploit geometric properties, utilize intermediate data, such as similarities computed at previous timestamps, and adjust trajectory granularity. These strategies efficiently eliminate irrelevant trajectory comparisons early in the search process while maintaining accuracy. Together, they optimize continuous searches and reduce computational costs notably. The contributions are summarized as follows.
@@
  • We develop a new index structure, the segment-based vessel trajectory index (SVTI), which partitions and organizes historical vessel trajectories to accelerate similarity searches.
@@
The rest of the paper is organized as follows. Section 2 presents the problem and solution overview. Section 3 covers the segment-based vessel trajectory index. Section 4 details the new similarity measure. Section 5 presents the algorithms and speed-up strategies for continuous similar vessel trajectory search. Section 6 reports the experimental results. Section 7 reviews related works. Section 8 concludes the paper and offers future directions.
2. Preliminaries
Section 2.1 presents key definitions. Section 2.2 defines the research problem. Section 2.3 gives an overview of the ACTIVE framework. Frequently used notation is listed in Table 1.
Table 1. Notation
@@ 2.1. Trajectory
Figure 2. ACTIVE framework
@@
In Figure 1, $\mathcal{T}=\{T_{1},T_{2},T_{3},T_{4},T_{5}\}$. We simplify $T_{i}$ to $T$ when the index $i$ is not relevant.
@@
Building on Definition 2.3, a trajectory $T_{i}$ can be represented as a sequence of segments $T_{i}=\langle s_{i}^{1},s_{i}^{2},\dots,s_{i}^{u}\rangle$, where $s_{i}^{u}$ is the $u^{\textit{th}}$ segment of $T_{i}$. For example, in Figure 1, trajectory $T_{1}$ consists of two segments, $s_{1}^{1}$ and $s_{1}^{2}$, thus $T_{1}=\langle s_{1}^{1},s_{1}^{2}\rangle$, where $s_{1}^{1}=\langle p_{1}^{1},p_{1}^{2},p_{1}^{3},p_{1}^{4}\rangle$ and $s_{1}^{2}=\langle p_{1}^{5},p_{1}^{6},p_{1}^{7},p_{1}^{8}\rangle$.
@@
In Figure 1, at time $t$, $o_{t}=(p_{o}^{6},\langle p_{o}^{1},p_{o}^{2},p_{o}^{3},p_{o}^{4},$ $p_{o}^{5},p_{o}^{6}\rangle)$, which means its current position is $p_{o}^{6}$ and its trajectory up to $t$ is $\langle p_{o}^{1},p_{o}^{2},p_{o}^{3},p_{o}^{4},$ $p_{o}^{5},p_{o}^{6}\rangle$.
@@
Continuing Example 2.5, the pivotal point $p_{4}^{x}$ of $T_{4}$ w.r.t. $o$ is $p_{4}^{5}$, as it is the closest to $p_{o}^{6}$. Consequently, we use $(T_{4}^{\prime},T_{4}^{L})$ to represent $T_{4}$, where $T_{4}^{\prime}=\langle p_{4}^{1},p_{4}^{2},p_{4}^{3},p_{4}^{4},p_{4}^{5}\rangle$ and $T_{4}^{L}=L(p_{4}^{5},p_{4}^{7})$ is the dotted line between $p_{4}^{5}$ and $p_{4}^{7}$.
@@
Continuing Example 2.7, given the range $r$ in the example, the range query $\mathcal{R}(p_{o}^{c},r)$ returns $\mathcal{T}_{r}=\{T_{1},T_{3},T_{4}\}$.
2.2. Problem Definition
Research Problem 0 (Continuous Similar Trajectory Search (CSTS)).
@@ 2.2. Problem Definition
Here, $\mathcal{D}(o,T)$ is the similarity measure OTRD (defined in Section 4.2). Continuing Example 2.9, given $k=2$, $\Theta_{t}=\{T_{1},T_{4}\}$.
2.3. Framework Overview
Figure 2 shows the three key modules in ACTIVE framework:
@@ 2.3. Framework Overview
3. Segment-based Trajectory Index
3.1. Index Structure
The segment-based vessel trajectory index (SVTI) is a spatial access method designed to index vessel trajectory data efficiently. Each trajectory is divided into segments and a R-Tree structure organizes these segments hierarchically, enabling fast querying by grouping spatially close segments. Each leaf node stores the Minimum Bounding Rectangle (MBR) of its segments and a set of tuples, such as ($i,U$), where $i$ is the trajectory ID, and $U$ is the set of segment IDs within $T_{i}$. The leaf nodes also link to detailed information, including the sequence of points forming each segment and the search path for preceding segments, enabling seamless segment-to-trajectory mapping. Non-leaf nodes store the MBR of the node itself and references to trajectory and segment IDs in their child nodes. This hierarchical design preserves spatial relationships and trajectory continuity across different tree levels, ensuring efficient indexing and querying.
@@ 3.1. Index Structure
Figure 3 illustrates an example of SVTI. The index includes two leaf nodes, $M_{11}$ and $M_{12}$. In $M_{11}$, four segments are stored, each corresponding to one of four trajectories: $T_{1}$, $T_{2}$, $T_{3}$, $T_{4}$. $M_{12}$ contains three segments, involving trajectories $T_{1}$, $T_{3}$, $T_{4}$. Each leaf node is linked to both the segment data and the connection information. For example, $M_{12}$ contains $s^{2}_{1}$, which is the second segment in $T_{1}$. Detailed segment information is associated with $s^{2}_{1}$, including its preceding segment $s^{1}_{1}$ and the path $\langle M_{0},M_{11}\rangle$ that leads to $s^{1}_{1}$. The two leaf nodes, $M_{11}$ and $M_{12}$, are then grouped into a higher-level node $M_{0}$, which serves as the root node in this case. $M_{0}$ manages a total of seven segments, involving four different trajectories.
3.2. Segment Partitioning

### v3

@@ Abstract
Publicly available vessel trajectory data is emitted continuously from the global AIS system. Continuous trajectory similarity search on this data has applications in, e.g., maritime navigation and safety. Existing proposals typically assume an offline setting and focus on finding similarities between complete trajectories. Such proposals are less effective when applied to online scenarios, where similarity comparisons must be performed continuously as new trajectory data arrives and trajectories evolve. We therefore propose a real-time continuous trajectory similarity search method for vessels (ACTIVE). We introduce a novel similarity measure, object-trajectory real-time distance, that emphasizes the anticipated future movement trends of vessels, enabling more predictive and forward-looking comparisons. Next, we propose an efficient continuous similar trajectory search (CSTS) algorithm together with a segment-based vessel trajectory index and a variety of search space pruning strategies that reduce unnecessary computations during the continuous similarity search, thereby further improving efficiency. Extensive experiments on two large real-world AIS datasets offer evidence that ACTIVE is capable of outperforming state-of-the-art methods considerably. ACTIVE significantly reduces index construction costs and index size while achieving a 70% reduction in terms of query time and a 60% increase in terms of hit rate.
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
I Introduction
@@
Automatic Identification System (AIS) data, which provides real-time information on vessel positions, speeds, and directions, offers a rich resource for analyzing vessel movements. With the growing availability of AIS data, understanding vessel trajectories has become essential for navigation safety [1] and environmental monitoring [2]. Trajectory similarity search [3] plays a key role in applications such as collision avoidance [4], route optimization [5], and movement pattern analysis [6].
@@
Given the dynamic nature of maritime environments, real-time continuous trajectory similarity search is particularly valuable. It enables timely and proactive decision-making by identifying similar movement patterns as new AIS data arrives, facilitating responsive navigation and operational safety. Unlike traditional methods, which rely on retrospective analysis of complete trajectories, real-time similarity search operates incrementally, processing incoming data without requiring an entire trajectory. Moreover, real-time search considers the future movement trends and the destination of a moving object. For instance, as illustrated in Fig. 1, for a moving vessel $o$, given its current trajectory $T_{o}$ and target destination $p_{o}^{d}$, we continuously search for similar historical trajectories, accounting for both past movements and the vessel’s future trend. Despite significant progress in real-time trajectory similarity search, existing studies [7, 8, 9, 10, 11, 12, 13, 14, 15, 16] face limitations, leading to the following challenges.
Figure 1: Example of trajectories at timestamp $t$.
@@
Challenge I: How to effectively measure trajectory similarity. Existing methods, such as Dynamic Time Warping [17, 18], Longest Common Subsequence [19], Fréchet distance [20, 21, 22], Hausdorff distance [23, 24], and Edit distance-based measures [25, 26, 27], are commonly used to identify similar trajectories by analyzing historical movement data. However, they rely heavily on static or full-path comparisons, making them of low utility for evolving trajectories that need to be analyzed incrementally as new data arrives. Next, they focus exclusively on historical movements, overlooking the importance of an object’s near-future movement and target destination. This limits their effectiveness in real-time applications, such as collision detection, traffic management, and route prediction, where anticipating future movements is crucial. For example, in Fig. 1, the most similar trajectory for $T_{o}$ is $T_{3}$ using Fréchet or Hausdorff distance. However, when considering the destination of the object, the most similar trajectory should be $T_{1}$.
@@
Challenge II: How to improve efficiency and scalability in continuous similarity trajectory search. Existing similar trajectory search methods [7, 8, 9, 10, 11, 12] are usually designed for searches using traditional similarity metrics intended for historical data, making them less effective for real-time applications. Moreover, they are designed for offline, batch processing on static datasets, requiring full re-computations when new data arrives. This is unsuitable in a setting with continuous updates to AIS data. Achieving efficient, real-time continuous similarity searches for vessel trajectories involves addressing several interconnected challenges. First, vessel trajectories are complex, often involving irregular and dynamic movements in open water. In contrast, on road networks, structured paths and intersections simplify trajectory analysis. Second, pruning irrelevant or low-similarity candidates with high accuracy is particularly complex in a dynamic environment where trajectory data changes continuously. Effective pruning methods must strike a balance between reducing candidate sets and maintaining the accuracy of similarity computations. Third, scaling a system and retaining fast response times when handling continuous queries over extended periods adds another layer of difficulty. The system must be capable of managing large-scale, real-time data streams while maintaining high accuracy and adapting to varying query loads. Together, these challenges call for new innovations to achieve a scalable and accurate trajectory search system.
Thus, we present ACTIVE, a novel framework for real-time continuous trajectory similarity search for vessels.
@@
To address Challenge I, we propose a novel similarity measure called object-trajectory real-time distance, which prioritizes the future trends of vessel movements and their target destinations while accounting also for past trajectory parts and the significance of individual trajectory points. This measure enables more accurate predictions of future vessel positions and behaviors, making it better suited for real-time decision-making in maritime navigation and traffic management.
@@
To address Challenge II, we design an efficient continuous similar trajectory search algorithm on a segment-based trajectory index with four speed-up strategies that exploit geometric properties, utilize intermediate data, such as similarities computed at previous timestamps, and adjust trajectory granularity. These strategies efficiently eliminate irrelevant trajectory comparisons early in the search process while maintaining accuracy. Together, they optimize continuous searches and reduce computational costs notably. The contributions are summarized as follows.
@@
The rest of the paper is organized as follows. Section II presents preliminaries. Section III details the new similarity measure. Section IV presents the algorithms for continuous similar vessel trajectory search. Section V reports the experimental results. Section VI reviews related works. Section VII concludes the paper and offers future directions.
II Preliminaries
Section II-A presents key definitions. Section II-B defines the research problem. Section II-C gives an overview of the ACTIVE. Frequently used notation is listed in Table I.
TABLE I: Notation.
@@ II-A Trajectory
Figure 2: ACTIVE framework.
@@
In Fig. 1, $\mathcal{T}=\{T_{1},T_{2},T_{3},T_{4},T_{5}\}$. We simplify $T_{i}$ to $T$ when the index $i$ is not relevant.
@@
Building on Definition 3, a trajectory $T_{i}$ can be represented as a sequence of segments $T_{i}=\langle s_{i}^{1},s_{i}^{2},\dots,s_{i}^{u}\rangle$, where $s_{i}^{u}$ is the $u^{\textit{th}}$ segment of $T_{i}$. For example, in Fig. 1, trajectory $T_{1}$ consists of two segments, $s_{1}^{1}$ and $s_{1}^{2}$, thus $T_{1}=\langle s_{1}^{1},s_{1}^{2}\rangle$, where $s_{1}^{1}=\langle p_{1}^{1},p_{1}^{2},p_{1}^{3},p_{1}^{4}\rangle$ and $s_{1}^{2}=\langle p_{1}^{5},p_{1}^{6},p_{1}^{7},p_{1}^{8}\rangle$.
@@
In Fig. 1, at time $t$, $o_{t}=(p_{o}^{6},\langle p_{o}^{1},p_{o}^{2},p_{o}^{3},p_{o}^{4},$ $p_{o}^{5},p_{o}^{6}\rangle)$, which means its current position is $p_{o}^{6}$ and its trajectory up to $t$ is $\langle p_{o}^{1},p_{o}^{2},p_{o}^{3},p_{o}^{4},$ $p_{o}^{5},p_{o}^{6}\rangle$.
@@
Continuing Example 1, the pivotal point $p_{4}^{x}$ of $T_{4}$ w.r.t. $o$ is $p_{4}^{5}$, as it is the closest to $p_{o}^{6}$. Consequently, we use $(T_{4}^{\prime},T_{4}^{L})$ to represent $T_{4}$, where $T_{4}^{\prime}=\langle p_{4}^{1},p_{4}^{2},p_{4}^{3},p_{4}^{4},p_{4}^{5}\rangle$ and $T_{4}^{L}=L(p_{4}^{5},p_{4}^{7})$ is the dotted line between $p_{4}^{5}$ and $p_{4}^{7}$.
@@
Continuing Example 2, given the range $r$ in the example, the range query $\mathcal{R}(p_{o}^{c},r)$ returns $\mathcal{T}_{r}=\{T_{1},T_{3},T_{4}\}$.
II-B Problem Definition
Research Problem (Continuous Similar Trajectory Search (CSTS)).
@@ II-B Problem Definition
Here, $\mathcal{D}(o,T)$ is the similarity measure OTRD (defined in Section III-B). Continuing Example 3, given $k=2$, $\Theta_{t}=\{T_{1},T_{4}\}$.
II-C Framework Overview
Fig. 2 shows the three key modules in ACTIVE framework:
@@ II-C Framework Overview
III Similarity Measure
III-A Case Analysis
Figure 3: Examples of gaps between existing similarity measures and real needs.
Case 1: Trend Neglect. Traditional trajectory similarity measures focus on minimizing geometric or temporal differences without considering the directional movement. In Fig. 3(a), the distance between trajectories $T_{1}$ and $T_{o}$ is computed to be the same as that between $T_{2}$ and $T_{o}$ under existing measures. However, in many real-world applications, such as predictive modeling or motion forecasting, the trajectory trend plays a critical role in determining similarity. For example, although $T_{2}$ is farther from $T_{o}$ at earlier timestamps, its trend more closely aligns with $T_{o}$’s current movement direction, suggesting that $T_{2}$ is more likely to converge with $T_{o}$ in the near future. This indicates that ignoring the trend can lead to suboptimal or misleading results, particularly in scenarios where the future behavior of trajectories is of primary interest.
Case 2: Segment Relevance. Conventional trajectory similarity measures assume that all segments or points along a trajectory contribute equally to the similarity computation. This assumption overlooks the fact that, in many downstream tasks, certain segments of a trajectory may carry more relevance than others. For example, in Fig. 3(b), given a trajectory $T_{o}$ and a set of trajectories $T_{1},T_{2}$}, only the segment highlighted in the red box is of primary interest for comparison. Ignoring the varying importance of trajectory segments can lead to suboptimal decision-making in critical scenarios.
Case 3: Target Point Neglect. In many practical applications, such as logistics and real-time navigation, the target destination of a moving object is of paramount importance. Current trajectory similarity measures, however, are predominantly backward-looking, focusing on historical data while disregarding the significance of the target and the trajectory’s future course. For instance, as shown in Fig. 3(c), while the historical paths of trajectories $T_{1}$ and $T_{2}$ are similar, $T_{1}$’ trend after the pivotal point is more aligned with the projected path of $T_{o}$ when the target destination is taken into consideration. Incorporating the target point into trajectory similarity measures would not only enhance the accuracy of real-time trajectory search but also enable more reliable predictions of future movements in applications such as autonomous navigation, shipping logistics, and traffic management.
Figure 4: Example of OTRD measure.
III-B Object-Trajectory Real-time Distance
Definition 7 (Object-Trajectory Real-time Distance).
Given a moving object $o$ at timestamp $t$ with $o_{t}=(p_{o}^{c},T_{o})$ and a trajectory $T_{i}=\langle p_{i}^{1},p_{i}^{2},\dots,p_{i}^{n}\rangle$, the Object-Trajectory Real-time Distance between $o$ and $T_{i}$ is defined as follows:
$\mathcal{D}(o,T_{i})=\alpha\cdot\mathcal{D}_{H}(T_{o},T^{\prime}_{i})+(1-\alpha)\cdot\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L}).$ (1)
[…]
$\mathcal{D}_{T}(p_{o}^{d},T_{i}^{L})=d(p_{o}^{d},L(p_{i}^{x},p_{i}^{n})).$ (3)
Here, $d(p_{o}^{*},p_{i}^{*})$ is the euclidean distance between $p_{o}^{*}$ and $p_{i}^{*}$, where $p_{o}^{*}$ and $p_{i}^{*}$ are last points in $T_{o}$ and $T^{\prime}_{i}$. $\theta\in[0,1]$ is the decay factor. $T^{\prime\prime}_{o}$ is the original $T_{o}$. $|T|$ is the number of points in trajectory $T$. $x=\text{argmin}_{0<j\leq n}d(p_{o}^{c},p_{i}^{j})$ is the ID of the pivotal point in $T_{i}$. $T_{o}^{*}$ and $T^{\prime}_{i}$$^{*}$ are the segments of $T_{o}$ and $T^{\prime}_{i}$ with the last point removed. The distance can be flexibly customized by a trade-off parameter $\alpha\in[0,1]$ according to specific application needs. The process of OTRD computation is detailed in Section IV-C.
@@ III-B Object-Trajectory Real-time Distance
Fig. 4 shows an example of OTRD. Given a trajectory $T_{1}=\langle(1,2),(3,3),(5,3),(6,5),(7,6),(9,7)\rangle$, a moving object $o$ at time $t$ with the target destination $p_{o}^{d}=(9,5)$ and $(p_{o}^{c},T_{o})$, where $p_{o}^{c}=(6,3)$ and $T_{o}=\langle(2,1),(4,2),(5,2),(6,3)\rangle$, the trade-off parameter $\alpha=0.9$, and the decay factor $\theta=0.9$, we have the pivotal point of $T_{1}$, $p_{1}^{x}=p_{1}^{3}=(5,3)$, and two parts of $T_{1}$, $T^{\prime}_{1}=\langle(1,2),(3,3),(5,3)\rangle$ and $T_{1}^{L}=L((5,3),(9,7))$. The HTD and TTD can be computed accordingly, i.e., $\mathcal{D}_{H}(T_{o},T^{\prime}_{1})=1.14$ and $\mathcal{D}_{T}(p_{o}^{d},T_{1}^{L})=1.41$. Therefore, the OTRD between $o$ and $T_{1}$ is $\mathcal{D}(o,T_{1})=0.9*1.14+0.1*1.41=1.167$.
IV Search Algorithm
In this section, we present an efficient search algorithm for CSTS, which operates on a segment-based trajectory index and incorporates four acceleration strategies.
IV-A Segment-based Trajectory Index
The segment-based vessel trajectory index (SVTI) is a spatial access method designed to index vessel trajectory data efficiently. Each trajectory is divided into segments and an R-Tree structure organizes these segments hierarchically, enabling fast querying by grouping spatially close segments. Each leaf node stores the Minimum Bounding Rectangle (MBR) of its segments and a set of tuples, such as ($i,U$), where $i$ is the trajectory ID, and $U$ is the set of segment IDs within $T_{i}$. The leaf nodes also link to detailed information, including the sequence of points forming each segment and the search path for preceding segments, enabling seamless segment-to-trajectory mapping. Non-leaf nodes store the MBR of the node itself and references to trajectory and segment IDs in their child nodes. This hierarchical design preserves spatial relationships and trajectory continuity across different tree levels, ensuring efficient indexing and querying.
Figure 5: An example of SVTI.
Example 5.
Fig. 5 illustrates an example of SVTI. The index includes two leaf nodes, $M_{11}$ and $M_{12}$. In $M_{11}$, four segments are stored, each corresponding to one trajectory: $T_{1}$, $T_{2}$, $T_{3}$, $T_{4}$. $M_{12}$ contains three segments, involving trajectories $T_{1}$, $T_{3}$, $T_{4}$. Each leaf node is linked to both the segment data and the connection information. For example, $M_{12}$ contains $s^{2}_{1}$, which is the second segment in $T_{1}$. Detailed segment information is associated with $s^{2}_{1}$, including its preceding segment $s^{1}_{1}$ and the path $\langle M_{0},M_{11}\rangle$ that leads to $s^{1}_{1}$. The two leaf nodes, $M_{11}$ and $M_{12}$, are then grouped into a higher-level node $M_{0}$, which serves as the root node in this case. $M_{0}$ manages a total of seven segments, involving four different trajectories.